# 0-supple.tex

### v1

@@ C.2 Proof of the Results for the Heterogeneous Setting
$\sup_{j}\left\lVert\widehat{V}_{j,t}\widehat{V}_{j,t}^{\top}-V_{j,t}V_{j,t}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{V}}\kappa_{0}\sigma_{t}}{\lambda_{\min}}\Big{(}1+\frac{\kappa^{2}\sqrt{r}\sigma_{s}}{\sqrt{\sigma^{2}_{s}+\sigma_{t}^{2}}}\Big{)}\right).$
C.3 Proof of the Results for Distributed Inference
Proof for Theorem B.1.
where $\mathfrak{E}^{(t)}_{j,\ell}$ is a remainder term defined by
$\zeta_{j,\ell,1}$ $:=T_{j}U_{-j}U_{-j}^{\top}Z^{\top}_{j,\ell},$
$\zeta_{j,\ell,3}$ $:=Z_{j,\ell}U_{-j}U_{-j}^{\top}Z_{j,\ell}^{\top},$
where
For each $j\in[J]$, $\ell\in[L]$, define a “locally-good” event:
$\quad\widetilde{E}_{j,\ell}(C)$ (42)
and a “globally-good” event
$\widetilde{E}(C):=\bigcap_{\ell=1}^{L}\bigcap_{j=1}^{J}E_{j,\ell}(C).$
By the proof of Lemma C.1 and Theorem 2.1, we have $\mathbb{P}[\widetilde{E}(C_{2})]\geq 1-C_{1}e^{-c_{1}p}$ for some constants $C_{1},c_{1},C_{2}$. By Lemma C.2, it holds that
$\quad-U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}+\mathfrak{R}^{(t)}_{j,\ell},$
where for $j\in[J]$ and $t=1,2$,
$\quad-U_{j\perp}U_{j\perp}^{\top}\zeta^{\top}_{j,\ell,1}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}(\zeta_{j,\ell,1}+\zeta^{\top}_{j,\ell,1})U_{j}\Lambda_{j}^{-2}U_{j}^{\top},$
and
$\mathfrak{S}^{(t)}_{j,\ell,3}$ $:=U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\zeta^{(t)}_{j,\ell,4}U_{j\perp}U_{j\perp}^{\top}+U_{j\perp}U_{j\perp}^{\top}\zeta^{(t)}_{j,\ell,4}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$ (47)
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j\perp}U_{j\perp}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j\perp}U_{j\perp}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j\perp}U_{j\perp}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}$
$\quad-U_{j\perp}U_{j\perp}^{\top}\mathfrak{E}^{(t)}_{j,\ell}U_{j}\Lambda_{j}^{-2}U_{j}^{\top}\left[\mathfrak{E}^{(t)}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top}\right]U_{j}\Lambda_{j}^{-2}U_{j}^{\top}+\mathfrak{R}^{(t)}_{j,\ell}.$
Then
By (30), it holds that
with probability at least $1-C_{1}e^{-c_{1}p}$.
Therefore, we obtain that
where
$\quad-U_{j}U_{j}^{\top}\overline{\mathfrak{S}_{j,1}}U_{j\perp}U_{j\perp}^{\top}\overline{\mathfrak{S}_{j,1}}U_{j}U_{j}^{\top}+\widetilde{\mathfrak{R}}_{j},$
Moreover,
Hence,
$\quad\zeta^{(2)}_{j,\ell,2}$ (53)
$\quad+\mathfrak{R}_{\zeta^{(2)}_{j,\ell,2}},$
where
$\mathfrak{M}_{j^{\prime}}:=\mathrm{tr}\Bigg{\{}\Lambda_{j}^{-4}G_{j}$ (55)
$W_{j^{\prime},\ell,a}:=U^{\top}_{j^{\prime}\perp}Z_{j^{\prime},\ell}U_{-j^{\prime}},$
$W_{j^{\prime},\ell,b}:=U^{\top}_{j\perp}Z_{j,\ell}(U_{1}\otimes\cdots\otimes U_{j^{\prime}-1}\otimes U_{j^{\prime}\perp}\otimes U_{j^{\prime}+1}\otimes\cdots\otimes U_{j-1}\otimes U_{j+1}\otimes\cdots\otimes U_{J}),$
$W_{j,\ell,c}:=U^{\top}_{j\perp}Z_{j,\ell}U_{-j},$
then $\mathfrak{M}_{j^{\prime}}$ can be simplified as
$\mathfrak{M}_{j^{\prime}}=\mathrm{tr}\left[\Lambda_{j}^{-4}G_{j}\frac{1}{L}\sum_{\ell=1}^{L}\left(I_{r_{1}}\otimes\cdots I_{r_{j^{\prime}-1}}\otimes\left(\Lambda^{-2}_{j^{\prime}}G_{j^{\prime}}W^{\top}_{j^{\prime},\ell,a}\right)\otimes I_{r_{j^{\prime}+1}}\cdots\otimes I_{r_{J}}\right)W^{\top}_{j^{\prime},\ell,b}\overline{W_{j,c}}G_{j}^{\top}\right],$
where $\overline{W_{j,c}}=(1/L)\sum_{\ell=1}^{L}W_{j,\ell,c}$.
Let $\mathrm{vec}$ denote the vectorization of a matrix. By assumption, $\mathrm{vec}(Z_{j^{\prime},\ell})\sim\mathcal{N}(0,\sigma^{2}I)$. Using the identity that $\mathrm{vec}(ABC)=(C^{\top}\otimes A)\mathrm{vec}(B)$ for all matrices $A,B,C$, we obtain that $\mathrm{vec}(W_{j^{\prime},\ell,a})\sim\mathcal{N}(0,\sigma^{2}I)$, i.e., the entries of $W_{j^{\prime},\ell,a}$ are i.i.d. $\mathcal{N}(0,\sigma^{2})$. Similarly, $\mathrm{vec}(W_{j^{\prime},\ell,b})\sim\mathcal{N}(0,\sigma^{2}I)$ and $\mathrm{vec}(W_{j,\ell,c})\sim\mathcal{N}(0,\sigma^{2}I)$. Furthermore, since $U_{j^{\prime}\perp}^{\top}U_{j^{\prime}}=0$, $\mathbb{E}[\mathrm{vec}(W_{j^{\prime},\ell,b})\mathrm{vec}(W_{j,\ell,c})^{\top}]=0$, which implies that $W_{j^{\prime},\ell,b}$ and $W_{j,\ell,c}$ are independent. Therefore, conditional on $\{W_{j^{\prime},\ell,a},W_{j,\ell,c}\}_{\ell=1}^{L}$,
$\quad\mathfrak{M}_{j^{\prime}}\mid\{W_{j^{\prime},\ell,a},W_{j,\ell,c}\}_{\ell=1}^{L}$ (56)
$\sim\mathcal{N}\left(0,\frac{\sigma^{2}}{L^{2}}\sum_{\ell=1}^{L}\left\lVert\overline{W_{j,c}}G^{\top}_{j}\Lambda_{j}^{-4}G_{j}\left(I_{r_{1}}\otimes\cdots I_{r_{j^{\prime}-1}}\otimes\left(\Lambda^{-2}_{j^{\prime}}G_{j^{\prime}}W^{\top}_{j^{\prime},\ell,a}\right)\otimes I_{r_{j^{\prime}+1}}\cdots\otimes I_{r_{J}}\right)\right\rVert^{2}_{F}\right).$
By the proof of Theorem 2.1, with probability at least $1-C_{1}e^{-c_{1}p}$, $\left\lVert W_{j^{\prime},\ell,a}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p}$ and $\left\lVert\overline{W_{j,c}}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p/L}$. Note that
$\mathrm{rank}\left[\overline{W_{j,c}}G^{\top}_{j}\Lambda_{j}^{-4}G_{j}\left(I_{r_{1}}\otimes\cdots I_{r_{j^{\prime}-1}}\otimes\left(\Lambda^{-2}_{j^{\prime}}G_{j^{\prime}}W^{\top}_{j^{\prime},\ell,a}\right)\otimes I_{r_{j^{\prime}+1}}\cdots\otimes I_{r_{J}}\right)\right]\leq\mathrm{rank}(G_{j})=r_{j}.$
$\left\lVert\overline{W_{j,c}}G^{\top}_{j}\Lambda_{j}^{-4}G_{j}\left(I_{r_{1}}\otimes\cdots I_{r_{j^{\prime}-1}}\otimes\left(\Lambda^{-2}_{j^{\prime}}G_{j^{\prime}}W^{\top}_{j^{\prime},\ell,a}\right)\otimes I_{r_{j^{\prime}+1}}\cdots\otimes I_{r_{J}}\right)\right\rVert_{F}\leq C_{2}^{2}pr^{1/2}L^{-1/2}\sigma^{2}\lambda_{\min}^{-3}.$
Hence,
$\mathrm{sd}(\mathfrak{M}_{j^{\prime}})$ $:=\sqrt{\frac{\sigma^{2}}{L^{2}}\sum_{\ell=1}^{L}\left\lVert\overline{W_{j,c}}G^{\top}_{j}\Lambda_{j}^{-4}G_{j}\left(I_{r_{1}}\otimes\cdots I_{r_{j^{\prime}-1}}\otimes\left(\Lambda^{-2}_{j^{\prime}}G_{j^{\prime}}W^{\top}_{j^{\prime},\ell,a}\right)\otimes I_{r_{j^{\prime}+1}}\cdots\otimes I_{r_{J}}\right)\right\rVert^{2}_{F}}$
$\lesssim L^{-1}pr^{1/2}\sigma^{3}\lambda_{\min}^{-3}.$
By (56), for any $\gamma>0$,
$\mathbb{P}\left(\left\lvert\mathfrak{M}_{j^{\prime}}\right\rvert\geq\mathrm{sd}(\mathfrak{M}_{j^{\prime}})\sqrt{2\gamma\log p}\quad\big{|}\quad\{W_{j^{\prime},\ell,a},W_{j,\ell,c}\}_{\ell=1}^{L}\right)\leq 2p^{-\gamma}.$
Therefore, with probability at least $1-C_{1}Le^{-c_{1}p}-2p^{-\gamma}$,
In conclusion, the second term in (52) can be bounded by
with probability at least $1-C_{1}JLe^{-c_{1}p}-2Jp^{-\gamma}$.
Next, we deal with the third term in (52), i.e.,
$\mathrm{tr}\left(\Lambda_{j}^{-4}U^{\top}_{j}\overline{\zeta_{j,3}}U_{j\perp}U_{j\perp}^{\top}\overline{Z_{j}}U_{-j}G_{j}^{\top}\right)=\frac{1}{L}\sum_{\ell=1}^{L}\mathrm{tr}\left[\Lambda_{j}^{-4}W_{j,\ell,d}W_{j,\ell,c}^{\top}\overline{W_{j,c}}G_{j}^{\top}\right]:=\widetilde{\mathfrak{M}_{0}},$
where $W_{j,\ell,d}:=U_{j}^{\top}Z_{j,\ell}U_{-j}$. Since $\mathrm{vec}([W_{j,\ell,d}W_{j,\ell,c}])=\mathrm{vec}([U_{j}U_{j\perp}]^{\top}Z_{j,\ell}U_{-j})\sim\mathcal{N}(0,\sigma^{2}I)$, we have
$\widetilde{\mathfrak{M}_{0}}\mid\{W_{j,\ell,c}\}_{\ell=1}^{L}\sim\mathcal{N}\left(0,\frac{\sigma^{2}}{L^{2}}\sum_{\ell=1}^{L}\left\lVert\Lambda_{j}^{-4}G_{j}\overline{W_{j,c}}^{\top}W_{j,\ell,c}\right\rVert_{\rm F}^{2}\right).$
Using that $\left\lVert W_{j,\ell,c}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p}$ and $\left\lVert\overline{W_{j,c}}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p/L}$ with probability at least $1-C_{1}e^{-c_{1}p}$, we have $\mathrm{sd}(\widetilde{\mathfrak{M}}_{0})\lesssim\sigma^{3}L^{-1}r^{1/2}p\lambda^{-3}_{\min}$. Therefore, with probability at least $1-C_{1}Le^{-c_{1}p}-2p^{-\gamma}$, we obtain that
For the fourth term in (52), note that
$=\frac{1}{L}\sum_{\ell=1}^{L}\mathrm{tr}\left[\Lambda_{j}^{-4}G_{j}W^{\top}_{j,\ell,d}\Lambda_{j}^{-2}G_{j}W_{j,\ell,c}^{\top}\overline{W_{j,c}}G_{j}^{\top}\right]+\frac{1}{L}\sum_{\ell=1}^{L}\mathrm{tr}\left[\Lambda_{j}^{-4}W_{j,\ell,d}G_{j}^{\top}\Lambda_{j}^{-2}G_{j}W_{j,\ell,c}^{\top}\overline{W_{j,c}}G_{j}^{\top}\right]$
$=:\widetilde{\mathfrak{M}}_{1}+\widetilde{\mathfrak{M}}_{2}.$
Combining (52), (57), (58) and (60) leads to that
if
or
C.4 Proof of Technical Lemmas
Proof of Lemma C.1.
@@ C.4 Proof of Technical Lemmas
By Lemma C.3, for any $t>0$, there exist $c,C$ such that
@@
$\mathbb{P}\left[\left\lVert Z_{j,\ell}U_{-j}\right\rVert_{2}\leq C_{2}\sigma\sqrt{p}\right]\geq 1-2e^{-c_{1}p},$ (63)
for some $c_{1},C_{2}>0$.
To show the third inequality in (19), we first define $\mathcal{B}^{p\times r}(X_{0},\varepsilon):=\{X\,|\,X\in\mathbb{R}^{p\times r},\left\lVert X-X_{0}\right\rVert_{2}\leq\varepsilon\}$. By Lemma 7 in Zhang and Xia (2018), for any $\varepsilon>0$, there exist $\mathcal{C}^{p,r}:=\big{\{}\overline{X}_{1},\dots,\overline{X}_{N}\big{\}}$ such that $\left\lVert\overline{X}_{i}\right\rVert_{2}\leq 1$ and $\mathcal{B}^{p\times r}(\bm{0},1)\subset\bigcup_{i=1}^{N}\mathcal{B}^{p\times r}\big{(}\overline{X}_{i},\varepsilon\big{)}$, with $N\leq(1+2/\varepsilon)^{pr}$. In particular, let $\varepsilon=1/(2J)$, and we have that, for any $(X_{1},\dots,X_{J})$ satisfying $X_{j}\in\mathcal{B}^{p_{j}\times r_{j}}(\bm{0},1)$, there exists $\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\mathcal{C}^{p_{2},r_{2}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}$ such that $\left\lVert X_{j}-\widetilde{X}_{j}\right\rVert_{2}\leq 1/(2J)$ for all $j$. Since
@@
$\leq\sum_{j^{\prime}\in[J]\setminus\{j\}}\left\lVert X_{j^{\prime}}-\widetilde{X}_{j^{\prime}}\right\rVert_{2}\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}$
$\leq\frac{1}{2}\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2},$
@@
$\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}$ (64)
@@
Therefore, it suffices to show that there exist $C_{1},c_{1},C_{2}$, s.t.
$\sup_{\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}}\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq C_{2}\sigma\sqrt{pr},$
with probability at least $1-C_{1}e^{-c_{1}p}$.
For a fixed tuple $\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}$, let $\Psi_{j,\ell}=Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}$. The rows of $\Psi_{j,\ell}$ are independent normal vectors $\mathcal{N}(0,\Sigma_{-j})$, where $\Sigma_{-j}=\sigma^{2}\big{(}\widetilde{X}_{1}^{\top}\widetilde{X}_{1}\otimes\widetilde{X}_{2}^{\top}\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}^{\top}\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}^{\top}\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}^{\top}\widetilde{X}_{J}\big{)}$, and hence $\Psi_{j,\ell}\Sigma_{-j}^{-1/2}$ has i.i.d. standard normal entries. Applying Lemma C.3 to $\Psi_{j,\ell}\Sigma_{-j}^{-1/2}$ and using that $\left\lVert\Psi_{j,\ell}\right\rVert_{2}\leq\left\lVert\Psi_{j,\ell}\Sigma_{-j}^{-1/2}\right\rVert_{2}\left\lVert\Sigma_{-j}^{-1/2}\right\rVert_{2}\leq\sigma\left\lVert\Psi_{j,\ell}\Sigma_{-j}^{-1/2}\right\rVert_{2}$ leads to
@@
$\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq C^{\prime}\sigma\sqrt{pr},$
with probability at least $1-2e^{-c\gamma^{\prime 2}pr}$, for any $\gamma^{\prime}>0$ and some $C^{\prime}>0$ depending on $\gamma^{\prime}$. This further leads to that
$\sup_{\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}}\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq C^{\prime}\sigma\sqrt{pr},$
with probability at least $1-2(4J+1)^{Jpr}e^{-c\gamma^{\prime 2}pr}$. By choosing $\gamma^{\prime}$ sufficiently large and combining the above inequality with (64), we finally obtain that there exist $C_{1},c_{1},C_{2}>0$, such that
$\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}\leq C_{2}\sigma\sqrt{pr},$ (65)
with probability at least $1-C_{1}e^{-c_{1}pr}$. Combining (62), (63), and (65) leads to $\mathbb{P}[E_{j,\ell}(C_{2})]\geq 1-C_{1}e^{-c_{1}p}$, implying that $\mathbb{P}[E(C_{2})]\geq 1-C_{1}Le^{-c_{1}p}$, for some constants $C_{1},c_{1},C_{2}>0$. ∎
Proof of Lemma C.2.

### v3

@@ D.3 Proof of the Results for the Knowledge Transfer Setting
$\sup_{j}\left\lVert\widehat{V}_{j,\ell}\widehat{V}_{j,\ell}^{\top}-V_{j,\ell}V_{j,\ell}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr_{V}}\kappa_{0}}{\lambda_{\min}}\Big{(}\sigma_{\ell}+\frac{\kappa\kappa_{0}\sqrt{r}}{\sqrt{\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}}+\frac{\kappa\kappa_{0}r\sqrt{L}}{\sqrt{p^{J-1}\sum_{\ell=1}^{L}\sigma_{\ell}^{-2}}}\Big{)}.$
D.4 Proof of Technical Lemmas
Proof of Lemma D.1.
@@ D.4 Proof of Technical Lemmas
By Lemma D.3, for any $t>0$, there exist $c,C$ such that
@@
$\mathbb{P}\left[\left\lVert Z_{j,\ell}U_{-j}\right\rVert_{2}\leq C_{Z}\sigma\sqrt{p}\right]\geq 1-2e^{-c_{Z}p},$ (60)
for some $c_{Z},C_{Z}>0$. The same argument leads to
$\mathbb{P}\left[\left\lVert\overline{Z}_{j}U_{-j}\right\rVert_{2}\leq C_{Z}\sigma\sqrt{p/L}\right]\geq 1-2e^{-c_{Z}p},$ (61)
since $\overline{Z}_{j}=\frac{1}{L}\sum_{\ell=1}^{L}Z_{j,\ell}\in\mathcal{N}(0,\sigma^{2}/L)$.
To show the third inequality in (24), we first define the ball $\mathcal{B}^{p\times r}(X_{0},\varepsilon):=\{X\,|\,X\in\mathbb{R}^{p\times r},\left\lVert X-X_{0}\right\rVert_{2}\leq\varepsilon\}$. By Lemma 7 in Zhang and Xia (2018), for any $\varepsilon>0$, there exist $\mathcal{C}^{p,r}:=\big{\{}\overline{X}_{1},\dots,\overline{X}_{N}\big{\}}$ such that $\left\lVert\overline{X}_{i}\right\rVert_{2}\leq 1$ and $\mathcal{B}^{p\times r}(\bm{0},1)\subset\bigcup_{i=1}^{N}\mathcal{B}^{p\times r}\big{(}\overline{X}_{i},\varepsilon\big{)}$, with $N\leq(1+2/\varepsilon)^{pr}$. In particular, let $\varepsilon=1/(2J)$, and we have that, for any $(X_{1},\dots,X_{J})$ satisfying $X_{j}\in\mathcal{B}^{p_{j}\times r_{j}}(\bm{0},1)$, there exists $\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\mathcal{C}^{p_{2},r_{2}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}$ such that $\left\lVert X_{j}-\widetilde{X}_{j}\right\rVert_{2}\leq 1/(2J)$ for all $j$. Since
@@
$\leq\sum_{j^{\prime}\in[J]\setminus\{j\}}\left\lVert X_{j^{\prime}}-\widetilde{X}_{j^{\prime}}\right\rVert_{2}\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,\forall j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}$
$\leq\frac{1}{2}\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,\forall j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2},$
@@
$\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,\forall j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}$ (62)
@@
Therefore, it suffices to show that there exist $\widetilde{C}_{1},\widetilde{c}_{1},\widetilde{C}_{2}$, s.t.
$\sup_{\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}}\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq\widetilde{C}_{2}\sigma\sqrt{pr},$
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$.
For a fixed tuple $\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}$, let $\Psi_{j,\ell}=Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}$. The rows of $\Psi_{j,\ell}$ are independent normal vectors $\mathcal{N}(0,\Sigma_{-j})$, where $\Sigma_{-j}=\sigma^{2}\big{(}\widetilde{X}_{1}^{\top}\widetilde{X}_{1}\otimes\widetilde{X}_{2}^{\top}\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}^{\top}\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}^{\top}\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}^{\top}\widetilde{X}_{J}\big{)}$, and hence $\Psi_{j,\ell}\Sigma_{-j}^{-1/2}$ has i.i.d. standard normal entries. Applying Lemma D.3 to $\Psi_{j,\ell}\Sigma_{-j}^{-1/2}$ and using that $\left\lVert\Psi_{j,\ell}\right\rVert_{2}\leq\left\lVert\Psi_{j,\ell}\Sigma_{-j}^{-1/2}\right\rVert_{2}\left\lVert\Sigma_{-j}^{-1/2}\right\rVert_{2}\leq\sigma\left\lVert\Psi_{j,\ell}\Sigma_{-j}^{-1/2}\right\rVert_{2}$ leads to
@@
$\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq C_{\gamma}^{\prime}\sigma\sqrt{pr},$
with probability at least $1-2e^{-c\gamma^{\prime 2}pr}$, for any $\gamma^{\prime}>0$ and some $C_{\gamma}^{\prime}>0$ depending on $\gamma^{\prime}$. This further leads to that
$\sup_{\big{(}\widetilde{X}_{1},\dots,\widetilde{X}_{J}\big{)}\in\mathcal{C}^{p_{1},r_{1}}\times\dots\times\mathcal{C}^{p_{J},r_{J}}}\left\lVert Z_{j,\ell}\big{(}\widetilde{X}_{1}\otimes\widetilde{X}_{2}\otimes\cdots\otimes\widetilde{X}_{j-1}\otimes\widetilde{X}_{j+1}\otimes\cdots\otimes\widetilde{X}_{J}\big{)}\right\rVert_{2}\leq C_{\gamma}^{\prime}\sigma\sqrt{pr},$
with probability at least $1-2(4J+1)^{Jpr}e^{-c\gamma^{\prime 2}pr}$. By choosing $\gamma^{\prime}$ sufficiently large and combining the above inequality with (62), we finally obtain that there exist $\widetilde{C}_{1},\widetilde{c}_{1},\widetilde{C}_{2}>0$, such that
$\sup_{\begin{subarray}{c}X_{j^{\prime}}\in\mathbb{R}^{p_{j^{\prime}}\times r_{j^{\prime}}}\\ \left\lVert X_{j^{\prime}}\right\rVert_{2}\leq 1,j^{\prime}\in[J]\setminus\{j\}\end{subarray}}\left\lVert Z_{j,\ell}\big{(}X_{1}\otimes X_{2}\otimes\cdots\otimes X_{j-1}\otimes X_{j+1}\otimes\cdots\otimes X_{J}\big{)}\right\rVert_{2}\leq 2\widetilde{C}_{2}\sigma\sqrt{pr},$ (63)
with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}pr}$. Combining (59), (60), (61), and (63) leads to that $\mathbb{P}[E(C^{\prime}_{2})]\geq 1-C^{\prime}_{1}Le^{-c^{\prime}_{1}p}$ for some constants $C^{\prime}_{1},c^{\prime}_{1},C^{\prime}_{2}>0$. ∎
Proof of Lemma D.2.